### v2

# Three-Dimensional MRI Reconstruction with 3D Gaussian Representations: Tackling the Undersampling Problem

Abstract
Three-Dimensional Gaussian representation (3DGS) has shown substantial promise in the field of computer vision, but remains unexplored in the field of magnetic resonance imaging (MRI). This study explores its potential for the reconstruction of isotropic resolution 3D MRI from undersampled k-space data. We introduce a novel framework termed 3D Gaussian MRI (3DGSMR), which employs 3D Gaussian distributions as an explicit representation for MR volumes. Experimental evaluations indicate that this method can effectively reconstruct voxelized MR images, achieving a quality on par with that of well-established 3D MRI reconstruction techniques found in the literature. Notably, the 3DGSMR scheme operates under a self-supervised framework, obviating the need for extensive training datasets or prior model training. This approach introduces significant innovations to the domain, notably the adaptation of 3DGS to MRI reconstruction and the novel application of the existing 3DGS methodology to decompose MR signals, which are presented in a complex-valued format.
Index Terms:
MRI reconstruction; Gaussian representation; Isotropic resolution
I Introduction
Magnetic Resonance Imaging (MRI) serves as a critical diagnostic tool in contemporary medicine, offering high-resolution images of the human body without employing ionizing radiation. Three-dimensional (3D) MRI enhances the functionality of conventional two-dimensional (2D) MRI by generating volumetric datasets that allow for manipulation and examination from various angles, thus providing a comprehensive view of human anatomy [1]. Unlike 2D MRI, which produces images in discrete slices, 3D MRI captures extensive data volumes in a single acquisition, often using isotropic voxels to achieve higher spatial resolution with better signal-to-noise ratio. This enhancement facilitates improved detail and clearer distinction of anatomical structures, especially in anatomically complex regions [2, 3]. However, 3D MRI is associated with several challenges, including prolonged acquisition time that increases the risk of patient movement, potentially compromising image quality and necessitating repeat scans. Additionally, the extended acquisition time can affect patient comfort, posing ongoing challenges in 3D MRI applications.
Parallel imaging [4, 5] is a technique used to decrease the acquisition time of 3D MRI, although it faces certain constraint such as struggling to produce images without artifacts when the undersampling factor is high. Recent developments in compressed sensing [6, 7, 8] offer alternatives to further reduce the acquisition time of 3D MRI while preserving image quality. Techniques such as SPIRiT [9] and P-LORAKS [10], which leverage iterative self-consistency and low-rank modeling of k-space data, have been shown to enhance the efficiency of parallel imaging, providing complementary methods to accelerate 3D MRI reconstruction. Additionally, restoring MRI from undersampled acquisitions uing deep-learning-based algorithms received much attention. For example, unrolled model is a technique that unrolls the iterations of optimization as layers in neural networks[11, 12]. The diffusion model is another scheme that treats MRI undersampling as a process of gradually adding Gaussian noise to simulate the degradation from full to undersampled data and learns to recover the fully-sampled distributions from undersampled ones with noises [13, 14, 15]. While these approaches provide promising results, training data for pretraining is required. To overcome the limitation of the need for training data for pretraining, patient-specific reconstruction schemes are proposed. These frameworks usually involve exploration in image space by forcing consistency in frequency space with certain constraints such as total-variation (TV) constraints or low-rank plus sparsity constraints. It was found that a randomly initialized convolutional neural network (CNN) can inherently capture low-level image statistics, serving as an effective implicit prior for solving inverse problems and such scheme is usually termed as deep image prior [16, 17]. This insight enables a randomly initialized convolutional decoder (ConvDecoder) [18] to model the MRI image space and recover the images, as the CNN architecture inherently imposes spatial constraints by capturing local pixel dependencies. Implicit neural representation (INR) is another state-of-the-art technique [19] for modeling 3D volumes by mapping continuous spatial coordinates to signal values using powerful spatial encoding.
Recently, 3D Gaussian representation (3DGS) has emerged as a notable alternative for INR within the realm of 3D modeling and rendering, offering a versatile and robust framework for efficiently rendering complex scenes. 3DGS models 3D volumes as collections of Gaussian points, enabling the representation of both geometric and visual attributes. The Gaussian mixture model (GMM) is conceptually similar to 3DGS but generates predictions as a combination of multiple Gaussian distributions. In GMM, the probability density is influenced by the distance of a pixel value from a learned mean value. In contrast, 3DGS calculates probability density based on spatial distance, and introduces an additional learnable parameter that multiplies with the probability density to produce the final output. Hou et al. [20] integrated GMM into their MRI reconstruction pipeline to remove the artifacts and noise produced in iterative optimization process.
While 3DGS has proven effective in computer vision and certain medical imaging applications, its potential in MRI remains largely unexplored. Several studies [21, 22, 23] have explored 3DGS for X-ray image reconstruction. However, these approaches generally fail to establish a direct connection between the Gaussian representation and volumetric (voxel-based) representations due to the lack of an explicit transformation between the two. To reconstruct the final voxelized volume, the splatting technique in [23] is used to generate cone beam computed tomography projections, followed by the FDK algorithm for volume reconstruction. The work of [24] was the first to bridge this gap by introducing a CUDA-accelerated approach that enables the efficient transformation of 3D Gaussian points into grid points within a 3D volume. However, The model still adopts a tile-based voxelization design to maintain efficiency. In this study, we aim to apply 3DGS to explore the undersampled MRI reconstruction problem within the context of isotropic resolution 3D MRI.
Recent studies have increasingly focused on training 3DGS in the frequency domain, proposing a progressive coarse-to-fine learning paradigm [25, 26, 27, 28, 29]. Specifically, FreGS [30] implements this approach by prioritizing low-to-high frequency components, leveraging low-pass and high-pass filters in Fourier space to enhance performance. These studies have inspired our work on using 3DGS for MRI reconstruction for high acceleration factors using the long-axis splitting strategy that we will introduce in the following section. The long-axis splitting strategy starts with only a limited number of Gaussian points to catch the coarse information, and by increasing the number of Gaussian points, we gradually recover the fine details in the images.
In this study, we introduce a novel approach employing 3D Gaussian representations, termed 3DGSMR, for reconstructing isotropic resolution 3D MRI from undersampled measurements. The 3DGSMR model is initialized with the inverse Fast Fourier Transform (iFFT) applied to undersampled k-space data, which serves as the foundation for subsequent optimizations aiming at detailed image recovery. The voxelized 3D MR volume is derived from the 3D Gaussian representation. To maintain data consistency, we implement data-consistency constraint in k-space to ensure that the Fourier transform of the voxelized volume aligns with the undersampled k-space data. 3DGSMR has demonstrated exceptional performance in the task of undersampled 3D MRI reconstruction. It effectively decomposes the complex-valued MRI data into 3DGS. Additionally, we have refined the densification strategies for 3DGS within this work, improving the speed of convergence, and enhancing the overall performance of the reconstruction for high acceleration factors.
The contributions and innovations of this research include: 1) The usage of a novel 3D MRI representation model. In this model, each voxel of the 3D MRI is represented by a combination of learnable 3D Gaussian points. This innovative approach facilitates a continuous representation of the image, enhancing the granularity and flexibility of the imaging process; 2) The expansion of the existing 3DGS framework to enable the decomposition of complex-valued 3D volumes. This extension broadens the applicability of 3DGS, allowing for more sophisticated manipulation and analysis of MRI data; 3) The refinement of the optimization of the original 3DGS to fit the isotropic resolution 3D MRI setting, and hence accelerates the speed of convergence for high acceleration factors.
II Method
II-A 3D MRI reconstruction from undersampled data
The main focus of this work is to recover the 3D MR image $\mathbf{X}\in\mathbb{C}^{3}$ from the undersampled multichannel k-space measurements. Mathematically speaking, the images are acquired by multichannel measurement in k-space:
$\mathbf{b}=\mathcal{A}(\mathbf{X})+\mathbf{n},$ (1)
where $\mathbf{n}$ is zero mean Gaussian noise matrix that corrupts the measurements during the acquisition, and $\mathcal{A}$ is the forward operator consisting of coil sensitivity maps, undersampling mask and Fast Fourier Transformation (FFT).
To reconstruct $\mathbf{X}$ from the undersampled measurements $\mathbf{b}$, the following minimization problem needs to be solved:
[…]
where $\mathcal{A}(\mathbf{X})-\mathbf{b}||^{2}$ is the data-consistency term and $\mathcal{R}(\mathbf{X})$ is a regularization term to penalize the solution. $\lambda$ is the balancing parameter that balances the data-consistency term and the regularization term.
II-B 3D Gaussian representation model for 3D MRI
In this work, we model each voxel in the 3D MRI volume $\mathbf{X}=[x_{j}]$ as the combination of a set of learnable 3D Gaussian points. Here, $\mathbf{x_{j}}\in\mathbb{C}$ is the element at spatial location $\mathbf{j}=(x_{j},y_{j},z_{j})$. With this, we can write
$x_{j}=\sum_{i=1}^{M}G_{i}^{3}(\mathbf{j}|\rho_{i},\mathbf{p}_{i},\mathbf{\Sigma}_{i}),$ (3)
where $G_{i}^{3}\in\mathbb{G}^{3}$ is a local 3D Gaussian representation component that defines a local Gaussian distribution density field. It is formulated as
$G_{i}^{3}(\mathbf{j}|\rho_{i},\mathbf{p}_{i},\mathbf{\Sigma}_{i})=\rho_{i}\cdot\exp\left(-\frac{1}{2}(\mathbf{j}-\mathbf{p}_{i})^{T}\mathbf{\Sigma_{i}}^{-1}(\mathbf{j}-\mathbf{p}_{i})\right),$ (4)
where $\rho_{i}\in\mathbb{C}$, $\mathbf{p}_{i}\in\mathbb{R}^{3}$, and $\mathbf{\Sigma_{i}}\in\mathbb{R}^{3\times 3}$ are learnable parameters representing the central density value, position of the center and covariance matrix in the Gaussian point, respectively. Note that $\rho_{i}$ is in the complex domain to match the complex-valued form of the MR images in our task. $\mathbf{p}_{i}$ defines the position of the Gaussian point cloud, while $\mathbf{\Sigma_{i}}$ defines the shape size and orientation. $\mathbf{\Sigma_{i}}$ can further be decomposed into rotation and scaling as
$\mathbf{\Sigma_{i}}=\mathbf{R}_{i}\mathbf{S}_{i}\mathbf{S}_{i}^{T}\mathbf{R}_{i}^{T},$
where $\mathbf{R}_{i}$ and $\mathbf{S}_{i}$ are the rotation and scaling matrix. The rotation matrix is constructed using quaternions with 4 parameters governing the rotation of the Gaussian point. These parameters collectively describe the 3D Gaussian representation derived from a normal 3D Gaussian distribution.
The optimization process of 3DGSMR can be viewed as exploring various combinations of Gaussian point clouds, accounting for different scales, rotations and complex values, to accurately model the MRI volume. This process is guided by loss functions, typically comprising L1 or L2 norms alongside some regularizations, which serve as constraints to direct the optimization efforts. Several factors influence the training process including the method of initialization, adaptive density control [31], and the implementation of constraints or regularization techniques.
II-C Initialization of 3D Gaussian representation
Initialization plays a vital role in the training of 3D Gaussian framework due to the vast range of diverse possible combinations of Gaussian point clouds. Usually, a coarse and primitive reconstruction result is used to guide the 3DGS initialization, from which the model can evolve. In the MRI setting, the iFFT of undersampled k-sapce data is used as the initialization in this work, as iFFT from undersampled k-sapce data is freely accessible in the setting of MRI reconstruction. In the iFFT reconstruction, we randomly sample $M$ grid points as the center of the initial Gaussian points. The quantity of initial Gaussian points plays a critical role in shaping the performance of the model. Then we set the scales of the initial Gaussian points as the average of the distances to the three nearest grid points and assume zero rotations. For the central density values of the Gaussian points for both the imaginary and real parts, they are initialized by scaling down with factor $k$ from the values of the sampled grid points.
II-D Adaptive control
Adaptive control functions as a mechanism for densifying Gaussian points when the current number of Gaussian points is insufficient to accurately represent the MR volume. It is triggered when the normalized gradients of the Gaussian centers exceed a set threshold $s$. In such cases, cloning or splitting is needed based on the current sizes of Gaussian points. If the sizes of the current Gaussian points exceed a threshold $s$, splitting will be activated; otherwise, cloning will be triggered. For the cloning, another Gaussian point of same shape and scale is added in the same position. The central density values are scaled to half. The newly cloned Gaussian points will be trained to fill the empty spaces that are under-reconstructed. Such process is illustrated in Figure 1.
Figure 1: Cloning method in adaptive control. The cloned Gaussian point will be further optimized to fit into unreconstructed region.
In terms of the splitting, the Gaussian point is split into two identical Gaussian points, each with size equals to the original Gaussian size scaled by a factor of 1/1.6, and the central values are scaled down by half. These setting are based on the design of primitive 3DGS [32]. Unlike cloning, the positions of the two newly split Gaussian points follow a normal distribution centered around the previous center position. This results in two split Gaussian points having different proportions of overlaps. This process is illustrated in Figure 2 (a). Deng et al. [28] indicated that excessive overlaps not only increase computational costs but also introduce blurring artifacts, which obscure fine details and hinder the accuracy of scene reconstruction. To overcome this problem, a long-axis splitting approach is proposed. This splitting approach removes cloning and splits the Gaussian points along the longest axis to half, while the other two axes are scaled by a factor of 0.85 of the original size. The density values of each child Gaussian point are scaled to 0.6 of original values. Based on [28], such setting ensures that the two Gaussian points after splitting do not overlap on the longest axis, ensuring consistent overall shape before and after the splitting. Such concept is depicted in Figure 2 (b). We show in this work that the long-axis splitting approach improves the reconstruction of high-frequency details and improves the reconstruction results for high acceleration factors.
Figure 2: (a) Original splitting method in adaptive control. (b) The long-axis splitting approach. A large Gaussian point splits into two small identical Gaussian points along the longest axis which are then further optimized to fit into unreconstructed region.
To improve the computational efficiency, the Gaussian points that the magnitude of the complex-valued central density values are too small while with large loss gradients are removed during the process of adaptive control.
II-E Voxelization
[…]
II-F Complexity analysis
Compared to voxel-based representations, 3DGS offers a more compact encoding in terms of spatial complexity. In the proposed model to represent MRI volume, each Gaussian point typically requires only twelve parameters to store (3 for center location, 3 for scaling, 4 for rotation quaternions, and 2 for real and imaginary part of the complex density value), and significantly fewer Gaussian points are needed to represent a volume than in voxelization. For instance, in our model, we limit the number of Gaussian points to no more than 400,000 for a MRI volume. This results in a total of $400,000\times 12=4,800,000$ parameters, which is similar to storing a $130\times 130\times 130\times 2=4,394,000$ image volume in a dense voxel grid. While is our setting and other real MRI applications, the image volume size is usually way larger than $130\times 130\times 130$, and the proposed framework provides a memory-efficient image representation.
Because of the introduction of tile-based design and CUDA kernel, the time complexity of 3DGS has been reduced significantly. Suppose there are $M$ voxels in the volume and $T$ tiles ($8\times 8\times 8$ in our setting), the average number of voxels in a single tile is $M/T$. Suppose there is an average of $N$ Gaussian points in each tile ($400,000/625=640$ in our setting), the time complexity for voxelization is M/TNM/T*N because computation on tiles is parallelized. Fourier transform is applied on the voxelized volume to compute loss after voxelization.
II-G Training
Figure 3: Training pipeline and components of 3DGSMR. (a) Overall pipeline that shows the data flow and gradient flow from initialization to the final output. The process includes initialization, voxelization, TV regularization, Fourier Transformation, and adaptive control. (b) shows the illustration of the voxelizer, which aggregates the sparse 3DGS at corresponding grid locations by performing a weighted sum, forming the voxelized volume. (c) depicts the original adaptive control that was proposed for 3DGS. In the cloning process, a Gaussian point is added at the same location, replicating the original one. During splitting, a single Gaussian point is divided into two separate Gaussian points, with their locations following normal distributions. Gaussians with small amplitude values are pruned and removed. (d) id the modified adaptive control that is designed to handle high acceleration factors in this work. Cloning is eliminated, and splitting is performed along the longest axis without causing overlap, preserving the integrity of the original shape.
Based on the above described optimization pattern for MRI reconstruction, the following loss function is solved for reconstructing the isotropic resolution 3D MRI:
$\mathcal{L}(\rho_{i},\mathbf{p}_{i},\mathbf{\Sigma_{i}})$ $=$ $||\mathcal{A}\left([\sum_{i=1}^{M}G_{i}^{3}(\mathbf{j}|\rho_{i},\mathbf{p}_{i},\mathbf{\Sigma}_{i})]\right)-\mathbf{b}||$
$+$ $\lambda\cdot TV\left(|\sum_{i=1}^{M}G_{i}^{3}(\mathbf{j}|\rho_{i},\mathbf{p}_{i},\mathbf{\Sigma}_{i})|\right).$
Note that the TV loss was applied to the magnitude of the 3D MR volumes. The overall training scheme is depicted in Figure 3. The voxelizer is responsible for converting 3D Gaussian representations into voxelized volumes and vice versa, enabling seamless transformation between 3DGS and voxelized representations.
In the setting of our training process, the maximum number of Gaussian points is set to 400 thousand (400k) to stop further densification. The densification is done in every 100 iterations.
III Experiments
III-A Dataset
This research study was conducted using data acquired from human subjects. Seven fully-sampled k-space sets acquired from seven healthy volunteers were used in this work to simulate undersampling scenario of different acceleration factors. The Institutional Review Board at the University of Texas Southwestern Medical Center approved the acquisition of the data, and written consents were obtained from all subjects. The data was acquired on a 1.5T scanner (Ingenia, Philips Healthcare, Best, Netherlands) with a 20-array dStream Head Neck coil. The datasets were acquired in a fully-sampled fashion using a single-shot T1-weighted inversion recovery turbo field echo (TFE) sequence. The sequence parameters include: field of view = 350 mm $\times$ 250 mm $\times$ 200 mm corresponding to Foot-Head direction $\times$ Anterior-Posterior direction $\times$ Right-Left direction, TE/TR = 3.4/7.2 ms, flip angle = 10, acquired voxel size = 0.9 mm $\times$ 0.9 mm $\times$ 0.9 mm, inversion time (TI) = 1122 ms, TFE factor = 303, and shot interval = 2300 ms. The acquisition time for each dataset is 6 minutes and 24 seconds.
We used an automatic algorithm implemented on the scanner to pre-select the best coils, that provide the best signal-to-noise-ratio in the region of interest. A PCA-based coil combination using singular value decomposition is used such that the approximation error is << 5%.
To simulate clinical scenarios, we applied 3D masks built based on stacked of 2D masks sampled from a 2D Gaussian distribution, and such masks enabled undersampling in phase-encoding direction and the readout direction is fully sampled. Furthermore, the masks have a fully sampled center calibration region to ensure better incorporation of the low-resolution regions of MRI images. To strike a balance between acceleration and performance, our experiments tested acceleration factors of 2, 4, 6, 8, 10, 12, 14, and 16 respectively.
III-B Implementation details
The proposed 3DGSMR scheme was implemented based on the PyTorch library and realized using a single NVIDIA A6000 graphic card that provides 48 GB of GPU memory. The GPU memory consumption was depended on number of Gaussian points. The model was initialized with a fixed number of Gaussian points based on the iFFT images obtained from the undersampled k-space. These initial Gaussian points were then trained based on the loss function and densified by cloning and splitting until the maximum allowable number of Gaussian points was reached. Three quantitative metrics, structural similarity index measure (SSIM) [35], peak signal-to-noise-ratio (PSNR) [35], and Learned Perceptual Image Patch Similarity (LPIPS) [36] were calculated to assess the reconstruction performance and determine the optimal stopping criteria for training. SSIM evaluates structural similarity between images, focusing on luminance, contrast, and texture. PSNR measures pixel-level differences, assessing noise and distortion, but lacks perceptual sensitivity. LPIPS captures perceptual similarity by comparing high-level features with deep neural networks, reflecting human visual perception of textures, shapes, and image quality.
The training process was progressive, continuing until the SSIM, PSNR, and LPIPS metrics reached a plateau, meaning that progressive training would bring little or even no more improvement.
IV Results
IV-A Ablation studies
In this section, we detail the process of the determination of the hyper-parameters in the proposed framework. The dataset from one subject was used to identify the optimal hyper-parameters in the proposed scheme. We then used the hyper-parameters to generate the experimental results for all the remaining datasets reported in this paper.
We first conducted the ablation study on the number of of initial Gaussian points $M$, scaling factor $k$, threshold $s$ between splitting and cloning, and the coefficient of the TV loss $\lambda$. The study was conducted on a single dataset with an acceleration factor of 8, and the training was performed with the same number of iterations. The results are shown in Table I. The highlighted parameters in the table are the ones that we picked in the proposed framework. In summary, we selected M=200M=200k, $k=0.1$, and $s=0.01$ to achieve an optimal solution that balances between performance and training efficiency.
[…]
$M=$ 50k 0.936 32.47 0.111 11min27s
$M=$ 100k 0.938 32.13 0.105 9min36s
$M=$ 200k 0.941 31.76 0.097 9min40s
$M=$ 400k 0.935 32.05 0.106 11min36s
$k=0.1$ 0.941 31.90 0.097 9min33s
$k=0.15$ 0.940 32.34 0.096 9min30s
$k=0.2$ 0.941 32.19 0.096 9min33s
$k=0.5$ 0.940 32.48 0.101 9min28s
$s=0.01$ 0.941 32.50 0.097 9min37s
$s=0.02$ 0.941 32.60 0.103 14min21s
$s=0.05$ 0.941 32.76 0.103 14min46s
$s=0.1$ 0.941 32.41 0.100 17min2s
$\lambda=0$ 0.939 31.68 0.097 9min36s
$\lambda=0.1$ 0.942 32.87 0.095 9min38s
$\lambda=0.5$ 0.941 32.21 0.097 9min37s
$\lambda=1$ 0.940 31.83 0.097 9min34s
$\lambda=100$ 0.941 32.06 0.097 9min31s
TABLE I: Ablation study results with optimal choices highlighted.
Figure 4: Visual comparison with SSIM, PSNR, and LPIPS of the reconstructed results from original optimization method versus new optimization method that introduces long-axis splitting and abandoning cloning along with initializing using only 500 Gaussian points.
[…]
original cloning & original splitting & $M=200000$ 0.911 30.28 0.131 9min36s
original cloning & original splitting & $M=500$ 0.804 29.02 0.268 24min3s
original cloning & long-axis splitting & $M=500$ 0.912 30.53 0.144 17min33s
original splitting & $M=500$ 0.794 28.95 0.276 24min35s
long-axis splitting & $M=200000$ 0.915 30.17 0.125 10min31s
long-axis splitting & $M=500$ 0.919 30.94 0.121 17min22s
TABLE II: Quantitative results for different adaptive control strategies on one subject with an acceleration factor of 10.
Figure 5: Illustration of the convergence progress. (a) Acceleration factor of 8. The proposed method reaches acceptable results after 200 iterations. In subsequent iterations, additional details are recovered, leading to improved image quality. (b) Acceleration factor of 10 with long-axis splitting. Long axis splitting helps recover fine details at higher acceleration factors.
IV-B Splitting strategy for high acceleration factors
While with the original splitting strategy shown in Figure 2 (a), we noticed that the reconstruction results degraded significantly with blurring artifacts when the acceleration factor is high (e.g. 10-fold or 12-fold undersampling). To improve the reconstruction performance, a different splitting strategy is studied, inspired by the work of [28]. Specifically, we modified the original adaptive control by abandoning cloning, and applying long-axis splitting as shown in Figure 2 (b), and initializing with less number of Gaussian points (e.g. $M=500$). We investigated such strategy based on one dataset with acceleration factor 10. The quantitative results obtained from different adaptive control strategies on this dataset are shown in Table II. It is suggested that initialization with 500 Gaussian points and using the long-axis splitting strategy and removing the cloning outperforms other combinations, indicating this modified optimization strategy has advantages in high acceleration factors. To further compare the reconstruction results for high acceleration factors, a visual comparisons between the original splitting strategy and the long-axis splitting strategy was shown in Figure 4. The results of the original splitting and cloning method show more noticeable distortions and blurring, whereas the optimization with long-axis splitting reduces these artifacts.
It is also worth mentioning here that for the low-to-mid acceleration factors (e.g. 2-fold, 4-fold, 6-fold, and 8-fold), different strategies used in Table II play no role on the reconstruction performance, indicating that the long-axis splitting strategy works only in the cases that the acceleration factors are high.
Figure 6: Showcase of the reconstruction results for acceleration factors 2, 3, 4, 6, 8, 10, 12, 14, and 16 from one dataset. The proposed framework demonstrates robust performance across different acceleration factors. As the acceleration factor continues to increase, the reconstructed images become more blurred, hindering the details in the images. Two regions in the selected coronal slice are zoomed in for better visual comparison.
IV-C Stopping criterion and reconstruction time
The experiments in this work are built on the premise that we begin with fully sampled images and manually apply undersampling masks to simulate undersampling scenarios. The reconstruction process is done without the usage of the fully sampled data anymore and only the quantitative results computation involves the usage of fully sampled data. During the reconstruction process, important metrics including SSIM, PSNR, and LPIPS were recorded for each iteration using TensorBoard. Since 3DGSMR can be trained iteratively and continuously, we implement a stopping criterion to halt training when both SSIM, PSNR, and LPIPS either plateaued as the reconstruction progressed.
In Figure 5, the intermediate reconstructions from one subject using acceleration factors 8 and 10 based on the original cloning and splitting and the long-axis splitting are shown, depicting how the image quality improves visually with the increase of the number of iterations.
It is observed that the time decreases as the acceleration factor increases. This can be attributed to the reduced constraints at higher acceleration factors, which allows for faster convergence to an optimal state. Note the reported time in ablation study included the computation of evaluation metrics for every iteration. Actual reconstruction time is halved without the evaluation of the quantitative metrics.
IV-D Effectiveness of the proposed framework
This section presents the reconstruction results across different acceleration factors. Figure 6 shows the reconstructions for different acceleration factors using the proposed framework as well as direct iFFT on one dataset. Two regions in the selected coronal slice containing gray matters and white matters are magnified for further visual comparison. As the acceleration factor increases, the white matter and gray matter regions from the iFFT reconstruction become increasingly blurred and noisy, making it difficult to distinguish the details. While the proposed 3DGSMR framework is able to resolve these artifacts and reconstruct the details in the images.
In Figure 6, quantitative results using SSIM, PSNR, and LPIPS are also present on each image. The iFFT results are used as the baseline to highlight the improvement of the proposed framework. The improvements from the quantitative perspective provided by the proposed 3DGSMR scheme demonstrate its effectiveness for isotropic resolution 3D MRI reconstruction.
IV-E Noise resistance characteristics of 3DGSMR
It is observed that 3DGSMR could effectively reduce noise in the reconstruction. In this section, we present results demonstrating the noise resistance capabilities of the proposed framework. We manually add noise sampled from Gaussian distribution to the acquired k-space to simulate the noise scenario. The reconstruction results are shown in Figure 7. It can be observed that the noise has been effectively reduced, while the texture details are well-preserved.
Figure 7: Visual comparison of the reconstructed images from noise-corrupted data, assessed using SSIM, PSNR, and LPIPS. The third image shows the recovery without undersampling, while the fourth image demonstrates recovery from 4-fold undersampling.
IV-F Comparison study
This section validates the effectiveness and robustness of the proposed method by exhibiting qualitative and visual comparison with other well-known and state-of-the-art isotropic resolution 3D MRI reconstruction methods introduced in the introduction section such as compressed-sensing with total variation regularization (CS-TV) [37], ConvDecoder[18], and INR [38]. The later two are deep-learning based reconstruction methods that perform in a subject-specific fashion which allows reconstruction from undersampled k-space data without any priors and pre-training. Reconstruction results are evaluated and compared using metrics SSIM, PSNR, and LPIPS throughout the dataset for different acceleration factors.
Figure 8: Qualitative and quantitative comparisons of the proposed 3DGSMR with CS-TV, ConvDecoder, and INR are presented for all the seven datasets, evaluating acceleration factors of 2, 4, 6, 8, 10, and 12 across SSIM, PSNR, and LPIPS metrics.
Figure 9: Visual comparison of the reconstruction results for CS-TV, ConvDecoder, INR, and proposed 3DGSMR on two subjects. (a) Visual comparison for acceleration factor of 8. (b) Visual comparison for acceleration factor of 10 with long-axis splitting. From the comparison, it can be found that the proposed scheme is able to recover the details with less distortions and artifacts.
In Figure 9, the visual comparison among different reconstruction methods are shown. Two axial slices from two different subjects together with two selected zoomed-in regions are present in the figure. The proposed method demonstrates superior performance visually as well as in SSIM and PSNR than the competing methods. Furthermore, the proposed method is able to restore the fine details in the images, closely resembling the ground truth. The improved image quality from the proposed reconstruction framework is also illustrated by the quantitative results, which are presented in Figure 8.
V Discussion
This study proposes an innovative approach for isotropic resolution 3D MRI reconstruction using 3DGS, leveraging its effectiveness to enable adaptive smoothing, interpolation, and precise spatial reconstruction from undersampled k-space data. This method has the potential to reduce the acquisition time for 3D MRI by highly undersample the k-space while maintaining reasonable image quality. Our experiments demonstrated that 3DGSMR can be effectively applied to isotropic resolution 3D MRI.
Our model proved memory efficiency. 3DGS can be seen as a form of compression of the original voxelized representation of the images. By using 3D Gaussian points, it reduces the amount of memory required for storing detailed volume data, while simultaneously enhancing the continuous representation between voxels. The voxelization pipeline with CUDA parallelization smooths the transformation between 3DGS and voxelized representation, making it more efficient in terms of both memory usage and speed.
The optimization scheme for 3DGS originated from [24] proved its robustness in reconstruction with low-to-mid acceleration factors (such as 2X, 4X, 6X, and 8X). By eliminating cloning and replacing the original splitting method with the new long-axis splitting, and initializing from sparse Gaussian points, the proposed 3DGSMR model delivers impressive results with high acceleration factors (10X and 12X). It is also observed that the results from 3DGSMR were less noisy and maintained continuity in content, demonstrating its ability to enhance image quality and improve visual appearance. We also demonstrated the model’s ability to resist noise disturbance in k-space data. The time spent on the reconstruction can be considered efficient if the process is halted once an acceptable standard is reached. Typically, it takes 600 iterations for low-to-mid acceleration factors and 1000 iterations for high acceleration factors that consume only few minutes to achieve clinically acceptable results. For acceleration factors greater than 12, the quality deteriorates significantly. For high acceleration factors, the original densification strategy introduces significant overlap in the densified Gaussian points, which amplifies the smoothing or blurring effect in the 3DGSMR reconstructions. This makes it harder for the model to converge towards small Gaussian points that capture fine details. The modified densification strategy that we proposed in this work uses only long-axis splitting without too much overlaps, allowing the Gaussian points to more easily split into smaller Gaussian points to recover the fined details. This approach facilitates the inference and recovery of the lost high-frequency components.
One of the advantages of 3DGSMR is that we could potentially apply the cutting-edged advancements of 3DGS into our proposed framework. Most recent works on 3DGS have focused on accelerating the optimization process or reducing aliasing artifacts [39, 29, 40, 41], while few have addressed memory efficiency or tackled the overfitting issues inherent in 3DGS [42, 43, 44]. Zhao et al. introduced parallelized distributed training pipelines for 3DGS, significantly reducing training time and alleviating high GPU memory demands. We plan to investigate these advancements and introduce them into the proposed 3DGSMR framework in the future work.
In this study, the 3D Gaussian points is mapped to grid points in k-space using Fourier Transform. The Gaussian distribution has a unique property where its Fourier transform remains a Gaussian distribution, meaning the frequency domain representation can be directly derived as combinations of Gaussian distributions from the 3DGS in image space. This approach is similar to the “Splatting” method in the original 3D Gaussian splatting [32]. Consequently, the Fourier transform of a combination of 3DGS samples results in a corresponding combination of 3DGS samples in the frequency domain. This principle holds true even with non-uniform sampling, providing a promising strategy to circumvent the computational overhead typically associated with extensive FFT, interpolation, or other non-uniform FFT techniques. We plan to investigate this further in the future work.
The study in this work focused on 3D isotropic MRI reconstruction, and limited in-house data prevented us from implementing and comparing other cutting-edge methods, which requires data for pre-training. As a result, we only compared the model with subject-specific methods that do not rely on pre-training or priors.
The proposed model also inherits some limitations of the original 3DGS framework, such as needle-like artifacts under high acceleration factors, which could be viewed in Figure 6. This suggests that a few Gaussian points may overfit without sufficient high resolution components in frequency domain for training. This issue becomes apparent in our study when mapping Gaussian points to grid points in the frequency domain. The number of Gaussian points in image space is not well balanced with the number of grid points in k-space, particularly in the high-resolution regions where the sampling is sparse. An alternative approach would be to also represent the frequency domain using Gaussian distributions. This would ensure that the number of points is balanced between the image and frequency domains. Furthermore, 3DGS faces challenges when processing low signal-to-noise ratio (SNR) images. The noise can be amplified during the Gaussian fitting process, leading to the introduction of artifacts or incorrect structures in the reconstructed image. The model may struggle to accurately differentiate between signal and noise when low SNR images are present, resulting in blurred details and inaccurate reconstructions, especially in high-frequency regions.
The convergence of 3DGS in medical image reconstruction remains an area with limited theoretical proof, particularly for clinical applications. While 3DGSMR shows promise in generating high-quality reconstructions for MRI, its ability to reliably converge to an ideal solution under various noise and data imperfections is not well-established. Furthermore, enhancing the model’s interpretability, especially in understanding how Gaussian distributions influence the final reconstruction, is crucial for clinical adoption. Mathematical convergence proofs and clearer explainability would be essential to ensure the model’s reliability and transparency in clinical decision-making.
VI Conclusion
In conclusion, this study has introduced and evaluated a novel approach, designated as 3DGSMR, which utilizes 3D Gaussian representations for the reconstruction of isotropic resolution 3D MRI from highly undersampled data. The 3DGSMR model has proven to be effective in managing the challenges associated with undersampled 3D MRI reconstruction, efficiently decomposing complex-valued MRI data using 3D Gaussian representations. Refinement were made in the densification strategies of the 3DGS, which have significantly accelerated the convergence process and improved the overall performance of the reconstructions for high acceleration factors. These advancements underscore the potential of 3DGSMR in enhancing the precision and efficiency of 3D MRI, thereby contributing valuable insights and tools to the field of medical imaging.